# Nested subspace learning with flags

### v1

@@ 2.3 Optimization on flag manifolds
There is a rich literature on optimization on smooth manifolds (Edelman et al., 1998; Absil et al., 2009; Boumal, 2023), and the particular case of flag manifolds has been notably addressed in Ye et al. (2022); Zhu and Shen (2024). Since flag manifolds can be represented as quotient spaces of Stiefel manifolds, which themselves are embedded in a Euclidean matrix space, one can develop some optimization algorithms without much difficulty. In this paper, we will use a steepest descent algorithm, which is drawn from several works (Chikuse, 2003; Nishimori et al., 2006; Ye et al., 2022; Zhu and Shen, 2024). Let $f\colon\operatorname{Fl}(p,q_{1:d})\to\mathbb{R}$ be a smooth function on a flag manifold, expressed in the Stiefel representation (e.g. $f(U_{1:d})=\sum_{k=1}^{d}\left\lVert{U_{k}}{U_{k}}^{\top}x\right\rVert$ for some $x\in\mathbb{R}^{p}$). Given $U_{1:d}\in\operatorname{St}(p,q)$, let $\operatorname{Grad}f(U_{1:d})=({\partial f}/{\partial U_{ij}})_{i,j=1}^{p,q}$ denote the (Euclidean) gradient of $f$. To “stay” on the manifold, one first computes the Riemannian gradient of $f$ at $U_{1:d}$, noted $\nabla f(U_{1:d})$. It can be thought of as a projection of the Euclidean gradient onto the tangent space and computed explicitly (Nishimori et al., 2006; Ye et al., 2022). Then, one moves in the opposite direction of $\nabla f(U_{1:d})$ with a so-called retraction, which is chosen to be the polar retraction of Zhu and Shen (2024, Eq. (49)), combined with a line-search. We iterate until convergence. The final steepest descent algorithm is described in Algorithm 1.
@@
     $[G_{1}|\dots|G_{d}]\leftarrow\operatorname{Grad}f(U_{1:d})$triangleright Euclidean gradient
     $\nabla\leftarrow\bigl{[}G_{k}-\bigl{(}U_{k}{U_{k}}^{\top}G_{k}+\sum_{l\neq k}U_{l}{G_{l}}^{\top}U_{k}\bigr{)}\bigr{]}_{k=1\dots d}$triangleright Riemannian gradient
     $U_{1:d}\leftarrow\operatorname{polar}(U_{1:d+1}-\alpha\nabla)$triangleright polar retraction + line search
@@
We can initialize Algorithm 1 randomly via Chikuse (2003, Theorem 1.5.5), or choose a specific flag depending on the application, as we will see in section 4.
@@
Many extensions of Algorithm 1 can be considered: conjugate gradient (Ye et al., 2022), Riemannian trust region (Absil et al., 2009) etc. We can also replace the polar retraction with a geodesic step (Ye et al., 2022) or other retractions (Zhu and Shen, 2024).
@@ 3 The flag trick in theory
In the remaining of the paper, we assume that the data has been already centered around a point of interest (e.g. its mean or geometric median), so that we are only interested in fitting linear subspaces and not affine ones. One could directly include the center in the optimization variables—in which case the geometry would be the one of affine Grassmannians (Lim et al., 2019) or affine flags (Pennec, 2018)—but we don’t do it in this work for conciseness.
@@ 3.1 From subspaces to flags of subspaces: the seminal example of PCA
PCA is known as the eigendecomposition of the sample covariance matrix. Originally, it can be formulated as the search for a low dimensional subspace that minimizes the unexplained variance (or maximizes the explained variance). Let $X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}$ be a data matrix with $n$ samples, let $\mathcal{S}\in\operatorname{Gr}(p,q)$ be a $q$-dimensional subspace, and let $\Pi_{\mathcal{S}}\in\mathbb{R}^{p\times p}$ be the orthogonal projection matrix onto $\mathcal{S}$. Then PCA consists in the following optimization problem on Grassmannians:
$\mathcal{S}_{q}^{*}=\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,\left\lVert X-\Pi_{\mathcal{S}}X\right\rVert_{F}^{2},$ (2)
where $\left\lVert M\right\rVert_{F}^{2}:=\operatorname{tr}\left(M^{\top}M\right)$ denotes the Frobenius norm. The solution to the optimization problem is the $q$-dimensional subspace spanned by the leading eigenvectors of the sample covariance matrix $S:=\frac{1}{n}XX^{\top}$, that we note $\mathcal{S}_{q}^{*}=\operatorname{Span}(v_{1},\dots,v_{q})$. It is unique when the sample eigenvalues $q$ and $q+1$ are distinct, which is almost sure when $q\leq\operatorname{rank}(S)$. We will assume to be in such a setting in the following for simplicity but it can be easily handled otherwise by “grouping” the repeated eigenvalues (cf. Szwagier and Pennec (2024, Theorem B.1)). In such a case, the principal subspaces are nested for increasing $q$, i.e., if $\mathcal{S}_{q}^{*}$ is the $q$-dimensional principal subspace, then for any $r>q$, one has $\mathcal{S}_{q}^{*}=\operatorname{Span}(v_{1},\dots,v_{q})\subset\operatorname{Span}(v_{1},\dots,v_{r})=\mathcal{S}_{r}^{*}$.
Another way of performing PCA is in a sequential manner (cf. footnote 1). We first estimate the 1D subspace $\mathcal{V}_{1}^{*}$ that minimizes the unexplained variance, then estimate the 1D subspace $\mathcal{V}_{2}^{*}$ that minimizes the unexplained variance while being orthogonal to the previous one, and so on and so forth. This gives the following constrained optimization problem on 1D Grassmannians:
$\mathcal{V}_{q}^{*}=\underset{\begin{subarray}{c}\mathcal{V}\in\operatorname{Gr}(p,1)\\ \mathcal{V}\perp\mathcal{V}_{q-1}\perp\dots\perp\mathcal{V}_{1}\end{subarray}}{\operatorname{argmin}}\,\left\lVert X-\Pi_{\mathcal{V}}X\right\rVert_{F}^{2}.$ (3)
This construction naturally yields a sequence of nested subspaces of increasing dimension—i.e. a flag of subspaces—best and best approximating the dataset:
@@
Those subspaces happen to be exactly the same as the ones obtained by solving the subspace learning optimization problem (2), although the way they are obtained (in a greedy manner) is different. This is generally not the case for other dimension reduction problems (for instance in robust subspace recovery, as it is raised in the final open questions of Lerman and Maunu (2018b)).
Hence, the optimal solution to the subspace learning formulation of PCA (2) is equivalent to the sequential formulation of PCA (3), and both yield a flag of subspaces best and best approximating the data. One can wonder if this result could be directly obtained by formulating an optimization problem on flag manifolds. The answer is yes, as first proven in Pennec (2018, Theorem 9) with an accumulated unexplained variance (AUV) technique, but there is not a unique way to do it. Motivated by the recent principal subspace analysis (Szwagier and Pennec, 2024), we propose in the following theorem a generic trick to formulate PCA as an optimization on flag manifolds.
@@
Let $X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}$ be a centered $p$-dimensional ($p\geq 2$) dataset with $n$ samples. Let $q_{1:d}:=(q_{1},q_{2},\cdots,q_{d})$ be a sequence of increasing dimensions such that $0<q_{1}<q_{2}<\cdots<q_{d}<p$. Let $S:=\frac{1}{n}XX^{\top}$ be the sample covariance matrix. Assume that it eigendecomposes as $S:=\sum_{j=1}^{p}\ell_{j}v_{j}{v_{j}}^{\top}$ where $\ell_{1}\geq\dots\geq\ell_{p}$ are the eigenvalues and $v_{1}\perp\dots\perp v_{p}$ are the associated eigenvectors. Then PCA can be reformulated as the following optimization problem on flag manifolds:
${\mathcal{S}_{1:d}^{*}}=\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\left\lVert X-\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}X\right\rVert_{F}^{2}.$ (5)
More precisely, one has ${\mathcal{S}_{1:d}^{*}}=\left(\operatorname{Span}(v_{1},\dots,v_{q_{1}}),\operatorname{Span}(v_{1},\dots,v_{q_{2}}),\dots,\operatorname{Span}(v_{1},\dots,v_{q_{d}})\right)$. The solution is unique if and only if $\ell_{q_{k}}\neq\ell_{q_{k+1}},\forall k\in[1,d]$.
@@
with $W=\sum_{k=1}^{d}(I_{p}-\Pi_{\mathcal{S}_{k}})$ and $S=\frac{1}{n}XX^{\top}$. Let $U_{1:d+1}:=[U_{1}|\dots|U_{d}|U_{d+1}]\in\mathcal{O}(p)$ be an orthogonal representative (cf. section 2) of the optimization variable $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$. Then one has $\leavevmode\nobreak\ {\Pi_{\mathcal{S}_{k}}=U_{1:d+1}\operatorname{diag}\left(I_{q_{k}},0_{p-q_{k}}\right){U_{1:d+1}}^{\top}}$. Therefore, one has $W=U_{1:d+1}\Lambda{U_{1:d+1}}^{\top}$, with $\leavevmode\nobreak\ {\Lambda=\operatorname{diag}\left(0\,I_{q_{1}},1\,I_{q_{2}-q_{1}},\dots,d\,I_{q_{d+1}-q_{d}}\right)}$. Hence, one has
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\left\lVert X-\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}X\right\rVert_{F}^{2}\Longleftrightarrow\underset{U\in\mathcal{O}(p)}{\operatorname{argmin}}\,\frac{n}{d^{2}}\operatorname{tr}\left(U\Lambda^{2}U^{\top}S\right).$ (9)
The latter problem is exactly the same as in Szwagier and Pennec (2024, Equation (19)), which solves maximum likelihood estimation for principal subspace analysis. Hence, one can conclude the proof on existence and uniqueness using Szwagier and Pennec (2024, Theorem B.1).
The key element of the proof of Theorem 4 is that averaging the nested projectors yields a hierarchical reweighting of the (mutually-orthogonal) principal subspaces. More precisely, the $k$-th principal subspace has weight $(k-1)^{2}$, and this reweighting enables to get a hierarchy of eigenspaces (Cunningham and Ghahramani, 2015; Pennec, 2018; Oftadeh et al., 2020). In the following, we note $\leavevmode\nobreak\ {\Pi_{\mathcal{S}_{1:d}}:=\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}}$ and call this symmetric matrix the average multilevel projector, which will be central in the extension of subspace methods into multilevel subspace methods.

### v2

@@ 2.3 Optimization on Flag Manifolds
There is a rich literature on optimization on smooth manifolds (edelman_geometry_1998; absil_optimization_2008; boumal_introduction_2023), and the particular case of flag manifolds has been notably addressed in ye_optimization_2022; zhu_practical_2024. Since flag manifolds can be represented as quotient spaces of Stiefel manifolds, which themselves are embedded in a Euclidean matrix space, one can develop some optimization algorithms without much difficulty. In this paper, we will use a steepest descent algorithm, which is drawn from several works (chikuse_statistics_2003; nishimori_riemannian_2006; ye_optimization_2022; zhu_practical_2024). Let $f\colon\operatorname{Fl}(p,q_{1:d})\to\mathbb{R}$ be a smooth function on a flag manifold, expressed in the Stiefel representation (e.g. $f(U_{1:d})=\sum_{k=1}^{d}\left\lVert{U_{k}}{U_{k}}^{\top}x\right\rVert$ for some $x\in\mathbb{R}^{p}$). Given $U_{1:d}\in\operatorname{St}(p,q)$, let $\operatorname{Grad}f(U_{1:d})=({\partial f}/{\partial U_{ij}})_{i,j=1}^{p,q}$ denote the (Euclidean) gradient of $f$. To “stay” on the manifold, one first computes the Riemannian gradient of $f$ at $U_{1:d}$, noted $\nabla f(U_{1:d})$. It can be thought of as a projection of the Euclidean gradient onto the tangent space and computed explicitly (nishimori_riemannian_2006; ye_optimization_2022). Then, one moves in the opposite direction of $\nabla f(U_{1:d})$ with a so-called retraction, which is chosen to be the polar retraction of zhu_practical_2024, combined with a line-search. We iterate until convergence. The final steepest descent algorithm is described in Algorithm 1.
@@
  $\begin{bmatrix}G_{1}&\cdots&G_{d}\end{bmatrix}\leftarrow\operatorname{Grad}f(U_{1:d})$ $\triangleright$ Euclidean gradient
  $\nabla\leftarrow\bigl[G_{k}-\bigl(U_{k}{U_{k}}^{\top}G_{k}+\sum_{l\neq k}U_{l}{G_{l}}^{\top}U_{k}\bigr)\bigr]_{k=1\dots d}$ $\triangleright$ Riemannian gradient
  $U_{1:d}\leftarrow\operatorname{polar}(U_{1:d}-\alpha\nabla)$ $\triangleright$ polar retraction + line search
@@
We can initialize Algorithm 1 “randomly” (under the invariant measure of james_normal_1954): we generate a random $p\times q$ matrix with normal entries and perform its polar decomposition; the orthogonal factor then follows a uniform distribution (chikuse_statistics_2003, Theorem 1.5.5). We can also initialize Algorithm 1 with a specific flag depending on the objective function, as we will see in Section 4. Such an initial flag can, for instance, be obtained via the eigendecomposition of a symmetric matrix related to the problem, or by refining the solution of the original subspace problem.
@@
Many extensions of Algorithm 1 can be considered: conjugate gradient (ye_optimization_2022), Riemannian trust region (absil_optimization_2008), etc. We can also replace the polar retraction with a geodesic step (ye_optimization_2022) or other retractions (zhu_practical_2024).
Remark 6 (Complexity Analysis)
The number of operations for the polar retraction is $O(pq^{2})+O(q^{3})$ (see zhu_practical_2024 and comparisons with other retractions therein). The number of operations for the Riemannian gradient (from the Euclidean gradient) is $O(pq^{2})$. The number of operations for the Euclidean gradient depends on the objective function $f$ (which itself depends on the dimension $p$, the signature $q_{1:d}$ and the number of samples $n$), and the overall complexity of Algorithm 1 also depends on the number of iterations (for both the main loop and the line search). Therefore, for completeness, we report the empirical running times of our steepest descent algorithm in Appendix D.
@@ 3 The Flag Trick in Theory
Remark 7 (Centering)
In the remaining of the paper, we assume that the data has been already centered around a point of interest (e.g. its mean or geometric median), so that we are only interested in fitting linear subspaces and not affine ones. One could directly include the center in the optimization variables—in which case the geometry would be the one of affine Grassmannians (lim_numerical_2019) or affine flags (pennec_barycentric_2018)—but we don’t do it in this work for conciseness.
@@ 3.1 From Subspaces to Flags of Subspaces: the Seminal Example of PCA
PCA is known as the eigendecomposition of the sample covariance matrix. Originally, it can be formulated as the search for a low dimensional subspace that minimizes the unexplained variance (or maximizes the explained variance). Let $X\coloneqq\begin{bmatrix}x_{1}&\cdots&x_{n}\end{bmatrix}\in\mathbb{R}^{p\times n}$ be a data matrix with $n$ samples, let $\mathcal{S}\in\operatorname{Gr}(p,q)$ be a $q$-dimensional subspace, and let $\Pi_{\mathcal{S}}\in\mathbb{R}^{p\times p}$ be the orthogonal projection matrix onto $\mathcal{S}$. Then PCA consists in the following optimization problem on Grassmannians:
$\mathcal{S}_{q}^{*}=\operatorname*{arg\,min}_{\mathcal{S}\in\operatorname{Gr}(p,q)}\left\lVert X-\Pi_{\mathcal{S}}X\right\rVert_{F}^{2},$ (2)
where $\left\lVert M\right\rVert_{F}^{2}\coloneqq\operatorname{tr}\left(M^{\top}M\right)$ denotes the Frobenius norm. The solution to the optimization problem is the $q$-dimensional subspace spanned by the leading eigenvectors of the sample covariance matrix $S\coloneqq\frac{1}{n}XX^{\top}$, that we note $\mathcal{S}_{q}^{*}=\operatorname{Span}(v_{1},\dots,v_{q})$. It is unique when the sample eigenvalues $q$ and $q+1$ are distinct, which is almost sure when $q\leq\operatorname{rank}(S)$. We will assume to be in such a setting in the following for simplicity but it can be easily handled otherwise by “grouping” the repeated eigenvalues (cf. szwagier_curse_2025). In such a case, the principal subspaces are nested for increasing $q$, i.e., if $\mathcal{S}_{q}^{*}$ is the $q$-dimensional principal subspace, then for any $r>q$, one has $\mathcal{S}_{q}^{*}=\operatorname{Span}(v_{1},\dots,v_{q})\subset\operatorname{Span}(v_{1},\dots,v_{r})=\mathcal{S}_{r}^{*}$.
Another way of performing PCA is in a sequential manner (cf. Remark 1). We first estimate the 1D subspace $\mathcal{V}_{1}^{*}$ that minimizes the unexplained variance, then estimate the 1D subspace $\mathcal{V}_{2}^{*}$ that minimizes the unexplained variance while being orthogonal to the previous one, and so on and so forth. This gives the following constrained optimization problem on 1D Grassmannians:
$\mathcal{V}_{q}^{*}=\operatorname*{arg\,min}_{\begin{subarray}{c}\mathcal{V}\in\operatorname{Gr}(p,1)\\ \mathcal{V}\perp\mathcal{V}_{q-1}\perp\dots\perp\mathcal{V}_{1}\end{subarray}}\left\lVert X-\Pi_{\mathcal{V}}X\right\rVert_{F}^{2}.$ (3)
This construction naturally yields a sequence of nested subspaces of increasing dimension—i.e. a flag of subspaces—better and better approximating the dataset:
@@
Those subspaces happen to be exactly the same as the ones obtained by solving the subspace learning optimization problem (2), although the way they are obtained (in a greedy manner) is different. This is generally not the case for other dimension reduction problems (for instance in robust subspace recovery, as it is raised in the final open questions of lerman_overview_2018).
Hence, the subspace learning formulation of PCA (2) is equivalent to the sequential formulation of PCA (3)—in terms of cumulative spans (4)—and both yield a flag of subspaces best and best approximating the data. One can wonder if this result could be directly obtained by formulating an optimization problem on flag manifolds. The answer is yes, as first proven in pennec_barycentric_2018 with an accumulated unexplained variance (AUV) technique, but there is not a unique way to do it. Motivated by the recent principal subspace analysis (szwagier_curse_2025), we propose in the following theorem a generic principle to formulate PCA as an optimization on flag manifolds.
@@
Let $X\coloneqq\begin{bmatrix}x_{1}&\cdots&x_{n}\end{bmatrix}\in\mathbb{R}^{p\times n}$ be a centered $p$-dimensional ($p\geq 2$) dataset with $n$ samples. Let $q_{1:d}\coloneqq(q_{1},q_{2},\dots,q_{d})$ be a sequence of increasing dimensions such that $0<q_{1}<q_{2}<\dots<q_{d}<p$. Let $S\coloneqq\frac{1}{n}XX^{\top}$ be the sample covariance matrix. Assume that it eigendecomposes as $S\coloneqq\sum_{j=1}^{p}\ell_{j}v_{j}{v_{j}}^{\top}$ where $\ell_{1}\geq\dots\geq\ell_{p}$ are the eigenvalues and $v_{1}\perp\dots\perp v_{p}$ are the associated eigenvectors. Then PCA can be reformulated as the following optimization problem on flag manifolds:
${\mathcal{S}_{1:d}^{*}}=\operatorname*{arg\,min}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\left\lVert X-\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}X\right\rVert_{F}^{2}.$ (5)
More precisely, one has ${\mathcal{S}_{1:d}^{*}}=\left(\operatorname{Span}(v_{1},\dots,v_{q_{1}}),\operatorname{Span}(v_{1},\dots,v_{q_{2}}),\dots,\operatorname{Span}(v_{1},\dots,v_{q_{d}})\right)$. The solution is unique if and only if $\ell_{q_{k}}\neq\ell_{q_{k}+1},\forall k\in\{1,\dots,d\}$.
@@
with $W=\sum_{k=1}^{d}(I_{p}-\Pi_{\mathcal{S}_{k}})$ and $S=\frac{1}{n}XX^{\top}$. Let $U_{1:d+1}\coloneqq\begin{bmatrix}U_{1}&\cdots&U_{d}&U_{d+1}\end{bmatrix}\in\mathcal{O}(p)$ be an orthogonal representative (cf. Section 2) of the optimization variable $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$. Then one has $\penalty 10000\ {\Pi_{\mathcal{S}_{k}}=U_{1:d+1}\operatorname{diag}\left(I_{q_{k}},0_{p-q_{k}}\right){U_{1:d+1}}^{\top}}$. Therefore, one has $W=U_{1:d+1}\Lambda{U_{1:d+1}}^{\top}$, with $\penalty 10000\ {\Lambda=\operatorname{diag}\left(0\,I_{q_{1}},1\,I_{q_{2}-q_{1}},\dots,d\,I_{q_{d+1}-q_{d}}\right)}$. Hence, one has
$\operatorname*{arg\,min}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\left\lVert X-\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}X\right\rVert_{F}^{2}\Longleftrightarrow\operatorname*{arg\,min}_{U\in\mathcal{O}(p)}\frac{n}{d^{2}}\operatorname{tr}\left(U\Lambda^{2}U^{\top}S\right).$ (9)
The latter problem is exactly the same as in szwagier_curse_2025, which solves maximum likelihood estimation for principal subspace analysis. Hence, one can conclude the proof on existence and uniqueness using szwagier_curse_2025.
The key element of the proof of Theorem 8 is that averaging the nested projectors yields a hierarchical reweighting of the (mutually-orthogonal) principal subspaces. More precisely, the $k$-th principal subspace has weight $(k-1)^{2}$, and this monotonic weighting enables to get a hierarchy of eigenspaces (absil_optimization_2008; cunningham_linear_2015; pennec_barycentric_2018; oftadeh_eliminating_2020). In the following, we note $\penalty 10000\ {\Pi_{\mathcal{S}_{1:d}}\coloneqq\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}}$ and call this symmetric matrix the average multilevel projector, which will be central in the extension of subspace methods into multilevel subspace methods.
Remark 9 (Weighted Projectors)
One can show that any convex combination of the nested projectors yields the same result. Indeed, let $\alpha_{1},\dots,\alpha_{d}$ be positive scalars. Let $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$ and $U_{1:d+1}\coloneqq\begin{bmatrix}U_{1}&U_{2}&\cdots&U_{d+1}\end{bmatrix}\in\mathcal{O}(p)$ be an orthogonal representative of $\mathcal{S}_{1:d}$ (cf. Section 2). Then, one has $\sum_{k=1}^{d}\alpha_{k}\Pi_{\mathcal{S}_{k}}=\sum_{k=1}^{d}\alpha_{k}\sum_{l=1}^{k}U_{l}{U_{l}}^{\top}=\sum_{l=1}^{d}\left(\sum_{k=l}^{d}\alpha_{k}\right)U_{l}{U_{l}}^{\top}$. Therefore, the weights in front of each (mutually-orthogonal) subspace $\operatorname{Span}(U_{l})$ are strictly decreasing, and the solutions of Theorem 8 are exactly the same. For criteria beyond PCA, the choice of weights might have an influence on the solution. In the following, we will by default consider uniform weights $\left(\frac{1}{d},\dots,\frac{1}{d}\right)$.